Noah drives 40 miles at a certain speed. Hitting serious traffic, he is forced to drive at a speed that is 60 mph slower. He drives at this new speed for 50 miles. If the entire trip takes 3 hours, what is his speed during each part of the trip?
Answer by mananth(16946)

Noah drives 40 miles --------------x mph speed.
50 miles ----------------(x-60) mph,
the entire trip takes 3 hours,
time part I + time part II=3
40/x + 50/(x-60)= 3
multiply by x(x-60)
40(x-60)+50x=x(x-60)*3
40x-2400+50x=3x^2-180x
3x^2-270x+2400=0
/3
x^2-90x+800=0
x^2-80x-10x+800=0
x(x-80)-10(x-80)=0
(x-80)(x-10)=0
x=80 OR 10
ignore 10

Noah drives 40 miles --------------80 mph
50 miles ----------------20 mph,